OpenStudy (anonymous):

Find the slope m and y intercept b of the line whose equation is given: 4x+3y=24

OpenStudy (anonymous):

do you know where to start?

OpenStudy (anonymous):

All I have is 3y=4x-24 and then I'm stuck

OpenStudy (anonymous):

divide both sides by 3. that means every term in both sides. you get y=(4/3)x-8 do you know what slope-intercept form is?

OpenStudy (anonymous):

I have a terrible teacher and am just going by the book. I know y=mx+b

OpenStudy (anonymous):

so the number with x is the slope and the number not with x is the y intercept if the equation is y=mx+b, m is the slope and b is the intercept

OpenStudy (anonymous):

So m=4/3 and b=8?

OpenStudy (anonymous):

b=-8 you have to include the - sign

OpenStudy (anonymous):

I get it! Thanks!